y eqauls x to the second power plus 6x plus 8
Answer by madya(1)

y = X2 + 6X + 8
y = x2 + 4x + 2x + 8 (by factorization)
y = x(x+4) +2(x+4) (taking common)
y = (x+2)(x+4)
THAT WAS A BUBBLE GUM QUESTION ACTUALLY...:-)